# Quantifying Very Extreme Precipitation and Temperature Using Huge Ensembles Generated by Machine Learning-based Climate Model Emulators

### v1

@@ Abstract
Weather extremes produce major impacts on society and ecosystems and are likely to change in likelihood and magnitude with climate change. However, very low probability events are hard to characterize statistically using observations or even climate model output because of short records/runs. For precipitation, consideration of such events arises in quantifying Probable Maximum Precipitation (PMP), namely estimating extreme precipitation magnitudes for designing and assessing critical infrastructure. A recent National Academies report on modernizing PMP estimation proposed using very large climate model-based ensembles to estimate extreme quantiles, possibly through machine learning-based ensemble boosting. Here we assess such an approach for the contiguous United States using a huge ensemble (10560 years) produced by a state-of-the-art emulator (ACE2) trained on ERA5 reanalysis. The results indicate that one can practically estimate very extreme precipitation and temperature quantiles, provided one uses appropriate statistical extreme value techniques. More specifically, the results provide evidence for (1) the use of threshold-exceedance methods with a sufficiently high threshold for reliable estimation (necessary for precipitation), (2) the robustness of results to variation in extremes by season and storm type, and (3) the sufficiency of the ensemble for well-constrained statistical uncertainty. Our results also show that the emulator produces extremes outside the range of the ERA5 training data. While this suggests that such emulators have potential for quantifying the climatology of extremes, we do not extensively investigate if this particular emulator is fit for purpose. Our focus is on how to use huge ensembles to estimate very extreme statistics, and we expect the results to be relevant for future, presumably improved emulators.
@@
Statistical analysis of 10560 years of simulated weather from a machine learning-based climate model emulator provides an approach to estimating the magnitude of very extreme precipitation and temperature events.
@@ 1 Introduction
Estimating the magnitude of very extreme weather events is essential for infrastructure design, risk assessment, and emergency preparedness. Impacts from extreme events can occur through flooding, heatwaves, drought, high winds, and other mechanisms. In this work, we are specifically interested in very low probability events, ones whose annual exceedance probability (AEP) could be orders of magnitude smaller than the reciprocal of the length of the observation record. For example, we may aim to estimate the magnitude of a 1-in-10000-year event (e.g., the 1-in-10000 AEP precipitation depth), or an event with even lower probability. (AEP values are also referred to as “return values” or “return levels”.) The short sample sizes available from observations or computationally-expensive climate models are inadequate to provide direct empirical estimates of such events.
An important use case for estimating such extreme events is the concept of PMP. PMP is a long-standing approach for quantifying very extreme precipitation for use in designing and assessing the safety of critical infrastructure such as dams and nuclear power plants. PMP has historically been defined as the maximum possible precipitation over a given area and time duration (NASEM, 2024). In the U.S., PMP has been estimated using an approach that combines a catalog of the most extreme observed events with storm transposition, a technique for trading space for time by judging where else an observed storm could plausibly occur. PMP estimates aim to account for various complicating factors such as orography and locational differences in climate. While motivated by PMP, we also consider temperature, and our results are relevant if one is interested in estimating magnitudes of very low probability events (i.e., extreme quantiles).
@@
A statistical extreme value analysis (EVA) typically aims to characterize the magnitude of extreme events, often beyond the range of the data. At EVA’s foundation is probability theory that characterizes the limiting distributions of random variables as they increase toward their upper limit (infinity if unbounded) (Coles, 2001). In practice, an EVA begins with the researcher taking a small subset of the most extreme data values from the full data (typically less than 5%). This subset is considered to provide the best information about the distribution’s tail, whose characterization is the critical component for estimating the magnitude of very low probability events. Two general approaches are widely used: analysis of block (e.g., annual) maxima or of threshold exceedances. Even though EVA was specifically designed to estimate far into a distribution’s tail, it can still be difficult to use EVA to obtain meaningful estimates of very low probability events. Simply put, it is difficult to take $\sim$100 years of data and say something relevant about a 1-in-10000 year event. When extrapolating far into the tail, EVA can provide a point estimate, but large uncertainty can limit its usefulness.
@@
Traditional climate and meteorological models are computationally expensive. Increased computing power has drastically increased our ability to run models at higher resolution and with better physical fidelity, and to create large ensembles to try to characterize uncertainty and variability (Flato et al., 2013; NASEM, 2024). Coupled with this has been a long history of developing stochastic weather generators (Ailliot et al., 2015), climate model emulators (Tebaldi et al., 2025), regional climate models (Giorgi, 2019) and other techniques to produce simulated weather comparable to (or at higher resolution than) that from models, but with less computation. Recent advances in machine learning (ML) have resulted in very fast emulators that perform better than traditional approaches (Bi et al., 2023; Lam et al., 2023; Price et al., 2025). At the time of NASEM (2024), ML-based emulators were used primarily at weather time scales for forecasting, but there are now very promising climate model emulators (Chapman et al., 2025; Cresswell-Clay et al., 2024; Watt-Meyer et al., 2025). This opens the possibility of generating huge ensembles and statistically analyzing the output to characterize extremes.
@@
While NASEM (2024) proposed a new definition and a climate model-based approach to estimating PMP, it expected that development of the new approach would occur over a period of years. The report anticipated that ensemble boosting might play a key role in increasing ensemble sizes. Indeed, less than two years after report publication, for this work we produced a huge ensemble with little difficulty. In particular, we use ACE2, for which extensive validation work has been done (Watt-Meyer et al., 2025) and for which code and input data are readily available. With this ensemble, we can investigate the statistical techniques one could use to analyze extremes using huge ensembles. Tying that investigation loosely to the NASEM recommendations, we note that the report anticipated that to avoid bias from inclusion of non-extreme precipitation values when fitting statistical distributions, one would want to use either very high thresholds for threshold exceedance analysis or very long blocks for block maxima-based analysis. A threshold-based approach can naturally exclude events from storm types or seasons that do not produce very extreme precipitation. Finally, the report presented sample size calculations showing that the sample size needed to reduce uncertainty sufficiently depended critically on the tail of the distribution. For tails that are not “too” heavy, one might need on the order of 1000 exceedances of the chosen threshold.
@@
  • Does a threshold exceedance-based analysis avoid the difficulties of storm typing or seasonal analysis?
@@
This allows us to assess the recommendations of NASEM (2024) empirically and give guidance for statistical analysis of huge ensembles beyond the PMP use case.
@@ 2.1 ACE2-ERA5 ensemble
We created an ensemble with a total of 10560 years of emulated weather, based on 12 initial conditions, run for 22 years of forcing data, and 40 repetitions (details in Appendix A). We aggregate to the daily scale by averaging the surface precipitation rate (before converting to precipitation per day) and taking the maximum 2m air temperature. For analysis we restrict to the grid cells contained in the contiguous U.S..
@@ 2.2 Statistical extreme value analysis
We use the climextRemes package (Paciorek, 2023) to carry out EVA separately for each grid cell, fitting both threshold exceedance and block maxima models using maximum likelihood. For block maxima analysis, we fit the generalized extreme value (GEV) distribution to annual maxima. For the threshold exceedance (peaks-over-threshold [POT]) analysis, we fit a point process representation of the generalized Pareto distribution (GPD) to observations exceeding a threshold (Smith, 1989; Coles, 2001, Section 7). Parametrizations for the GEV and GPD are given in Appendix A and correspond to Coles (2001). GPDs are fit to a series of increasing thresholds, corresponding to the upper 0.001 to 0.00001 quantiles of daily observations, spaced on a regular grid of 10 quantiles on the log scale. These correspond to selecting the largest 3857, 2313, 1387, 831, 499, 299, 180, 108, 65, and 39 observations. Note that our largest sample size still uses fewer than one observation per year, while our smallest sample size still requires extrapolation for 1-in-1000 year AEP values. AEP values are estimated using the parameter point estimates, with AEP value uncertainty calculated using the statistical delta method. For probabilities that are low but not too low, the size of our ensemble also allows us to assess EVA results by comparing AEP value estimates to empirical quantiles (e.g., Huang et al., 2016).
@@
Of particular interest will be the shape parameter $\xi$, which appears in both the GEV and GPD and which determines the upper tail’s behavior. If $\xi>0$ the upper tail is unbounded and heavy, if $\xi=0$ the upper tail is unbounded but light, and if $\xi<0$ the tail is bounded. Precipitation is often found to have a heavy tail (NASEM, 2024), while temperature is often found to be bounded (e.g., Philip et al., 2022; Bercos-Hickey et al., 2022). We also investigate a bias-variance trade-off inherent to extremes. Bias arises because the GEV or GPD become the correct distribution only as the block size increases to infinity or as the threshold approaches the distribution’s upper limit. As any data set is finite, increasing the block size or threshold decreases the number of retained observations, thereby increasing the uncertainty in estimating parameters or quantities based on the parameters. The practitioner must decide how to balance bias and variance, and estimates can be sensitive to block length or threshold (e.g., Ben Alaya et al., 2020).
@@ 2.3 Seasonality and storm types
A similar issue arises with storm types (NASEM, 2024). Impactful events could arise from different weather phenomena; for example, extreme precipitation could arise from tropical cyclones or strong convective storms. Thus, one might consider the overall precipitation distribution to be best represented statistically as a mixture of distributions over the different storm types, or, as an approximation, as a mixture over different seasons, to the extent that the storm types in a season are a subset of the types seen through the year. To account for the mixture, one could stratify the extreme data by type and perform separate analyses. Doing so fits data that are more homogeneous, better characterizing the distribution’s tail, which is the key to extrapolation. The challenge with this approach is that this requires data labeled by storm type, and AEP value estimation also requires estimation of the occurrence frequency of each storm type. Alternatively, one could ignore storm types, thereby essentially aggregating extreme behavior across storm types. The challenge with this approach is that aggregating over storm types carries the risk of bias when estimating the magnitude of very low probability events, especially if these most extreme events predominantly arise from a single storm type but the threshold exceedances include (presumably generally less extreme) values from other types. To estimate PMP, NASEM (2024) recommended a threshold exceedance approach applied to data unlabeled by storm type and advocated setting the threshold high enough to exclude most, if not all, of the data arising from storm types other than most extreme.
We explore the effect of fitting season-specific (DJF, MAM, JJA, SON) models under two approaches: (1) using the same thresholds as discussed above for each season (but excluding the two highest thresholds as these produce too-small sample sizes at the seasonal level) and (2) using the same number of exceedances in each season as in the full-year analysis. The former uses the same threshold and therefore smaller sample sizes (in some cases zero observations such as for temperature in winter in many locations), while the latter uses lower (or possibly equal for a season in a location in which all extremes occur in that season) thresholds and larger sample sizes.
@@ 3.1 Precipitation
Figure 1: Maximum daily precipitation from emulator for different numbers of simulated years compared to maximum in ERA5 (1940-2022). One-to-one lines are in red.
@@
Figure 2: Emulator-based extreme precipitation climatology: (a) 1-in-1000 year AEP depth estimates, (b) 1-in-100000 year AEP depth estimates, (c) shape parameter estimates, with corresponding standard error estimates (d), (e), and (f), respectively.
Next we assess the performance of the GEV and POT fitting approaches for estimation of AEP depths corresponding to shorter time periods, 1000 and 10000 years, that are within the range of our ensemble size. By investigating shorter periods, we can estimate AEP depths directly as empirical quantiles of the samples in the ensemble and use them to assess the performance of EVA-based AEP depth estimates. Fig. 3 shows that when the POT threshold is sufficiently high (e.g., using sample sizes of 499 and fewer), the AEP depth estimates are very similar to the empirical quantiles. In contrast, for lower thresholds (larger sample sizes), there appears to be a modest upward bias. In stark contrast, GEV estimation based on annual maxima appears to be severely upwardly-biased, particularly for grid cells with more extreme precipitation. The fact that the POT-based AEP depths are well-aligned with the empirical quantile-based estimates for shorter time periods provides support for using the POT-based AEP depth estimates for longer time periods (such as 100000 and one million years). Block maxima with blocks of a year are clearly insufficient for good estimation. One could explore using longer blocks (e.g., 5, 10 or 20 years), but in our exploration here we focus on POT methods and sensitivity to the threshold.
@@
We next explore the tail behavior of precipitation, in relation to the choice of threshold. Fig. 4 (top row) shows that shape parameter estimates systematically decrease (tending to move from above the 1:1 line to below it) as one increases the threshold (decreases the sample size of exceedances), with variability also increasing, particularly dramatically with small sample sizes. Fig. 12 shows that the number of positive shape parameter estimates decreases substantially, with the distributions shifting to the left (as well as becoming more variable with smaller sample sizes). Correspondingly, AEP depth estimates decrease as the threshold increases (Fig. 4, bottom row). While the AEP depth estimates do decrease systematically, for the larger threshold values, they largely stabilize. From an implementation perspective, this suggests that if one chooses a sufficiently high threshold (such as corresponding to $n=499$), one can reasonably estimate AEP depths with limited bias. From a conceptual perspective, the decrease in shape parameter estimates suggests that previous results showing positive estimates for precipitation could be driven at least in part by the (necessary) use of low thresholds, potentially causing bias by including exceedances not representative of the far tail of the distribution. Of course this emulator-produced output could fundamentally differ from actual precipitation.
@@
The analyses above ignore seasonality and storm types. Fig. 5 shows that if one fits season-specific POT models and then determines the overall AEP depth as the largest amongst the four seasons, the results are very similar to ignoring seasonality, provided the threshold is sufficiently high. This is consistent with extreme value theory, which requires a sufficiently high threshold to remove the influence of values from the bulk of the distribution. In the climate context, this removes smaller extremes corresponding to storm types (or seasons) that produce less extreme precipitation and naturally includes only the storm type(s) producing very extreme precipitation. From a practical perspective, being able to ignore storm type/seasonality is helpful, particularly avoiding the need to categorize events into types.
Figure 5: Comparison of 1-in-100000 year AEP depth estimates based on analysis of full-year data to maxima of seasonally-stratified estimates for each grid cell. Results are shown for the second approach discussed in the methods; results for the first approach are similar (Fig. 13). One-to-one lines are in red.
Finally, we turn from bias in estimation to focus on uncertainty, characterizing whether an ensemble of size 10560 is able to sufficiently reduce uncertainty in AEP depth estimates for practical use. As discussed in NASEM (2024), uncertainty is driven primarily by the sample size, by the tail behavior, and by the time duration. Fig. 6 shows the relative uncertainty (the standard error divided by the AEP depth estimate) for three durations when using $n=499$ exceedances. Except when considering the longest duration in combination with the most positive shape parameter estimates, relative uncertainty is generally less than 15%, corresponding to confidence intervals whose length is plus or minus 30% of the value of the estimate (e.g., an estimate of 20 cm with a confidence interval of (14, 26)). While the standard errors are approximate (see Methods), our purpose is to approximately characterize uncertainty without delving into the details of alternative methods (e.g., the bootstrap) for estimating uncertainty.
@@
For many applications/locations, a 10560-member ensemble may be sufficient. However, if not, it is feasible now, and will become more so, to produce even larger ensembles, while cautioning that relative statistical uncertainty decreases as the square root of the sample size.

### v2

@@ Abstract
Weather extremes produce major impacts on society and ecosystems and are likely to change in likelihood and magnitude with climate change. However, very low probability events are hard to characterize statistically using observations or even climate model output because of short records/runs. For precipitation, consideration of such events arises in quantifying Probable Maximum Precipitation (PMP), namely estimating extreme precipitation magnitudes for designing and assessing critical infrastructure. A recent National Academies report on modernizing PMP estimation proposed using very large climate model-based ensembles to estimate extreme quantiles, possibly through machine learning-based ensemble boosting. Here we assess statistical aspects of such an approach for the contiguous United States using a huge ensemble (10560 years) produced by a state-of-the-art emulator (ACE2) trained on ERA5 reanalysis. The results indicate that one can practically estimate very extreme precipitation and temperature quantiles, provided one uses appropriate statistical extreme value techniques. More specifically, the results provide evidence for (1) the use of threshold-exceedance methods with a sufficiently high threshold (necessary for precipitation) for reliable estimation, (2) the robustness of results to variation in extremes by season and storm type, and (3) the sufficiency of the ensemble for well-constrained statistical uncertainty. Our results also show that the emulator produces extremes outside the range of the ERA5 training data. While encouraging for emulators’ potential use for quantifying the climatology of extremes, more investigation is needed to assess whether emulators are fit for this purpose. Our focus is on how to use huge ensembles to estimate very extreme statistics; we expect the results to be relevant for future improved emulators.
@@
Weather extremes produce major impacts on society and ecosystems and are likely to change in likelihood and magnitude with climate change. Machine learning now provides methods to simulate weather for many hypothetical years at much lower computational expense than traditional models. This could provide the opportunity to estimate the magnitude of very low probability events using statistical methods applied to the simulation output, for use in planning, risk assessment, and adaptation. Our work finds that when used appropriately, statistical methods designed to analyze extremes can reliably estimate the magnitude of extremely high precipitation and temperature events in the United States, under the critical assumption that the machine learning method is a good mimic of real world weather.
[…]
Statistical analysis of 10560 years of simulated weather from a machine learning-based climate model emulator supports an approach to estimating the magnitude of very extreme precipitation and temperature events.
@@ 1 Introduction
Estimating the magnitude of very extreme weather events is essential for infrastructure design, risk assessment, and emergency preparedness. Impacts from extreme events can occur through flooding, heatwaves, drought, high winds, and other mechanisms. In this work, we are specifically interested in very low probability events, ones whose annual exceedance probability (AEP) could be orders of magnitude smaller than the reciprocal of the length of the observation record. For example, we may aim to estimate the magnitude of a 1-in-10000-year event (e.g., the 1-in-10000-year AEP precipitation depth), or an event with even lower probability. (AEP values are also referred to as “return values” or “return levels”.) The short sample sizes available from observations or computationally-expensive climate models are inadequate to provide direct location-specific empirical estimates of such events.
An important use case for estimating such extreme events is PMP, a long-standing concept for quantifying very extreme precipitation for use in designing and assessing the safety of critical infrastructure such as dams and nuclear power plants. PMP has historically been defined as the maximum possible precipitation over a given area and time duration (NASEM, 2024). In the U.S., PMP has been estimated using an approach that combines a catalog of the most extreme observed events with storm transposition (a technique for trading space for time by judging where else an observed storm could plausibly occur), moisture maximization, and orographic adjustment. While motivated by PMP, we also consider temperature.
@@
A statistical extreme value analysis (EVA) typically aims to characterize the magnitude of extreme events, often beyond the range of the data. At EVA’s foundation is probability theory that characterizes the limiting distributions of random variables as they increase toward their upper limit (infinity if unbounded) (Coles, 2001). In practice, an EVA begins with the researcher taking a small subset of the most extreme data values from the full data (typically much less than 5%). This subset is considered to provide the best information about the distribution’s tail, whose characterization is the critical component for estimating the magnitude of very low probability events. Two general approaches are widely used: analysis of block (e.g., annual) maxima or of threshold exceedances. Even though EVA was specifically designed to estimate far into a distribution’s tail, it can still be difficult to use EVA to obtain meaningful estimates of very low probability events. Simply put, it is difficult to take $\sim$100 years of data and say something relevant about a 1-in-10000 year event. When extrapolating far into the tail, EVA can provide a statistical estimate, but large uncertainty can limit its usefulness.
@@
Traditional climate and meteorological models are computationally expensive. Increased computing power has drastically increased our ability to run models at higher resolution and with better physical fidelity, and to create large ensembles to try to characterize uncertainty and variability (Flato et al., 2013; NASEM, 2024). Coupled with this has been a long history of developing stochastic weather generators (Ailliot et al., 2015), climate model emulators (Tebaldi et al., 2025), regional climate models (Giorgi, 2019) and other techniques to produce simulated weather comparable to (or at higher resolution than) that from models, but with less computation. Recent advances in machine learning (ML) have resulted in very fast emulators that perform better than traditional approaches (Bi et al., 2023; Lam et al., 2023; Price et al., 2025). At the time of NASEM (2024), ML-based emulators were used primarily at weather time scales for forecasting, but there are now promising climate model emulators (Chapman et al., 2025; Cresswell-Clay et al., 2024; Watt-Meyer et al., 2025). This opens the possibility of generating huge ensembles and statistically analyzing the output to characterize extremes.
@@
While NASEM (2024) proposed a new definition and a climate model-based approach for estimating PMP, it expected that development of the new approach would occur over a period of years. The report anticipated that ensemble boosting might play a key role in increasing ensemble sizes. Indeed, less than two years after report publication, for this work we produced a huge ensemble with little difficulty. In particular, we use ACE2, for which extensive validation work has been done (Watt-Meyer et al., 2025) and for which code and input data are readily available. With this ensemble, we can investigate the statistical techniques one could use to analyze extremes using huge ensembles. Tying that investigation loosely to the NASEM recommendations, we note that the report anticipated that to avoid bias from inclusion of non-extreme precipitation values when fitting statistical distributions, one would want to use either very high thresholds for threshold exceedance analysis or very long blocks for block maxima-based analysis. A threshold-based approach can naturally exclude events from storm types or seasons that do not produce very extreme precipitation. Finally, the report presented sample size calculations showing that the sample size needed to reduce uncertainty sufficiently depended critically on the tail of the distribution. For tails that are not “too” heavy, one might need on the order of 1000 exceedances of the chosen threshold.
@@
  • Can an appropriate threshold exceedance-based analysis avoid the difficulties of storm typing or seasonal analysis?
@@
Using an ML-based emulator trained on reanalysis provides a testbed to assess the recommendations of NASEM (2024) empirically and give guidance for statistical analysis of huge ensembles beyond the PMP use case. For real applications in the future, there are key advantages to using suitably-trained ML-based emulators: large sample sizes, the potential to produce realistic space-time fields, and the potential to produce out-of-sample events by training on the observed evolution of the earth system.
@@ 2.1 ACE2-ERA5 ensemble
We created an ensemble with a total of 10560 years of emulated weather, based on 12 initial conditions, run for 22 years of forcing data, and 40 repetitions (details in Appendix A). We aggregate to the daily scale by averaging the surface precipitation rate (before converting to precipitation per day) and taking the maximum 2m air temperature. For analysis we restrict to the grid cells contained in the contiguous U.S.
@@ 2.2 Statistical extreme value analysis
We use the climextRemes package (Paciorek, 2023) to carry out EVA separately for each grid cell, fitting both threshold exceedance and block maxima models using maximum likelihood. For block maxima analysis, we fit the generalized extreme value (GEV) distribution to annual maxima. For the threshold exceedance (peaks-over-threshold [POT]) analysis, we fit a point process representation of the generalized Pareto distribution (GPD) to observations exceeding a threshold (Smith, 1989; Coles, 2001, Section 7). Parametrizations for the GEV and GPD are given in Appendix A and correspond to Coles (2001). GPDs are fit to a series of increasing thresholds, corresponding to the upper 0.001 to 0.00001 quantiles of daily observations, spaced on a regular grid of 10 quantiles on the log scale. These correspond to selecting the largest 3857, 2313, 1387, 831, 499, 299, 180, 108, 65, and 39 observations. AEP values are estimated using the parameter point estimates, with AEP value uncertainty calculated using the statistical delta method on the MLE-based information matrix. For probabilities that are low but not too low, the size of our ensemble also allows us to assess EVA results by comparing AEP value estimates to empirical quantiles (e.g., Huang et al., 2016).
@@
Of particular interest will be the shape parameter $\xi$, which appears in both the GEV and GPD and which determines the upper tail’s behavior. If $\xi>0$ the upper tail is unbounded and heavy, if $\xi=0$ the upper tail is unbounded but light, and if $\xi<0$ the tail is bounded. Precipitation is often found to have a heavy tail (NASEM, 2024), while temperature is often found to be bounded (e.g., Philip et al., 2022; Bercos-Hickey et al., 2022). We also investigate a bias-variance trade-off inherent to extremes. Bias arises because the GEV or GPD become the correct distribution only as the block size increases to infinity or as the threshold approaches the distribution’s upper limit. As any data set is finite, increasing the block size or threshold decreases the number of retained observations, thereby increasing the uncertainty in estimating parameters or quantiles. The practitioner must decide how to balance bias and variance, and estimates can be sensitive to block length or threshold (e.g., Ben Alaya et al., 2020).
@@ 2.3 Seasonality and storm types
A similar issue arises with storm types (NASEM, 2024). Impactful events could arise from different weather phenomena; for example, extreme precipitation could arise from tropical cyclones or strong convective storms. Thus, one might consider the overall precipitation distribution to be best represented statistically as a mixture of distributions over the different storm types, or, as an approximation, as a mixture over different seasons, to the extent that the storm types in a season are a subset of the types seen through the year. To account for the mixture, one could stratify the extreme data by type and perform separate analyses. Doing so fits data that are more homogeneous, better characterizing the distribution’s tail, which is the key to extrapolation. The challenge with this approach is that this requires data labeled by storm type, and AEP estimation also requires estimation of the occurrence frequency of each storm type. Alternatively, one could ignore storm types, thereby essentially aggregating extreme behavior across storm types. The challenge with this approach is that aggregating over storm types carries the risk of bias when estimating the magnitude of very low probability events, especially if these most extreme events predominantly arise from a single storm type but the threshold exceedances include (presumably generally less extreme) values from other types. To estimate PMP, NASEM (2024) recommended a threshold exceedance approach applied to data unlabeled by storm type and advocated setting the threshold high enough to exclude most, if not all, of the data arising from storm types other than most extreme.
We explore the effect of fitting season-specific (DJF, MAM, JJA, SON) models under two approaches: (1) using the same thresholds as discussed above for each season (but excluding the two highest thresholds as these produce too-small sample sizes at the seasonal level) and (2) using the same number of exceedances in each season as in the full-year analysis. The former uses the same threshold and therefore smaller sample sizes (in some cases zero observations such as for temperature in winter in many locations), while the latter uses lower (or equal when all the annual exceedances occur in a single season) thresholds and larger sample sizes.
@@ 3.1 Precipitation
Figure 1: Maximum daily precipitation for each grid cell over the specified time period from emulator (for different numbers of simulated years) compared to maximum in ERA5 (1940-2022). One-to-one lines are in red.
@@
Figure 2: Emulator-based extreme precipitation climatology: (a) 1-in-1000 year AEP depth (cm) estimates, (b) 1-in-100000 year AEP depth (cm) estimates, (c) shape parameter estimates, with corresponding standard error estimates (d), (e), and (f), respectively.
Next we assess the performance of the GEV and POT fitting approaches for estimation of AEP depths corresponding to shorter time periods, 1000 and 10000 years, that are within the range of our ensemble size. By investigating shorter periods, we can estimate AEP depths directly as empirical quantiles of the samples in the ensemble and use them to assess the performance of EVA-based AEP depth estimates. Fig. 3 shows that when the POT threshold is sufficiently high (e.g., using sample sizes of 499 and fewer), the AEP depth estimates are very similar to the empirical quantiles. In contrast, for lower thresholds (larger sample sizes), there appears to be a modest upward bias. In stark contrast, GEV estimation based on annual maxima (the most common choice in practice) appears to be severely upwardly-biased, particularly for grid cells with more extreme precipitation. The fact that POT-based AEP depths are well-aligned with empirical quantile-based estimates for shorter time periods provides support for using POT-based AEP depth estimates for longer time periods (such as 100000 and one million years). Block maxima with blocks of a year are clearly insufficient for good estimation, as is also the case for POT estimation using the same sample size (results not shown). Use of longer blocks (e.g., 5, 10, 20, … years) for GEV analysis produces results (not shown) equivalent to those seen here for POT, but here we focus on POT methods and sensitivity to the threshold.
@@
We next explore the tail behavior of precipitation, in relation to the choice of threshold. Fig. 4 (top row) shows that shape parameter estimates systematically decrease (tending to move from above the 1:1 line to below it) as one increases the threshold (decreases the sample size of exceedances), with variability also increasing, particularly dramatically with small sample sizes. Fig. 12 shows that the number of positive shape parameter estimates decreases substantially, with the distributions shifting to the left (as well as becoming more variable with smaller sample sizes). Correspondingly, AEP depth estimates decrease as the threshold increases (Fig. 4, bottom row). While the AEP depth estimates do decrease systematically, for the larger threshold values, they largely stabilize. From an implementation perspective, this suggests that if one chooses a sufficiently high threshold (such as corresponding to $n=499$), one can reasonably estimate AEP depths with limited bias. From a conceptual perspective, the decrease in shape parameter estimates suggests that previous results showing positive estimates for precipitation could be driven at least in part by the (necessary) use of low thresholds, potentially causing bias by including exceedances not representative of the far tail of the distribution. Of course this emulator-produced output could fundamentally differ from actual precipitation, and the decrease in shape parameter estimates could be a result of the emulator output having a lighter tail than reality.
@@
The analyses above ignore seasonality and storm types. Fig. 5 shows that if one fits season-specific POT models and then determines the overall AEP depth as the value at which the sum of exceedance probabilities across the four seasons equals the desired probability, the results are very similar to ignoring seasonality, provided the threshold is sufficiently high. This is consistent with extreme value theory, which requires a sufficiently high threshold to remove the influence of values from the bulk of the distribution. In the climate context, this removes smaller extremes corresponding to storm types (or seasons) that produce less extreme precipitation and naturally includes only the storm type(s) producing very extreme precipitation. From a practical perspective, being able to avoid storm typing or the need to analyze seasons individually can be a helpful simplification, particularly avoiding the subjectiveness of and uncertainty from categorizing events into types.
Figure 5: Comparison of 1-in-100000 year AEP depth estimates based on analysis of full-year data to aggregation over seasonally-stratified estimates for each grid cell. Results are shown for the second approach discussed in the methods; results for the first approach are similar (Fig. 13). One-to-one lines are in red.
Finally, we turn from bias in estimation to focus on statistical uncertainty, characterizing whether an ensemble of size 10560 is able to sufficiently reduce uncertainty in AEP depth estimates for practical use. As discussed in NASEM (2024), uncertainty is driven primarily by the sample size, the tail behavior, and the exceedance probability of interest. Fig. 6 shows the relative uncertainty (the standard error divided by the AEP depth estimate) for three time periods when using $n=499$ exceedances. Except when considering the longest period in combination with the most positive shape parameter estimates, relative uncertainty is generally less than 15%, corresponding to confidence intervals whose length is plus or minus 30% of the value of the estimate (e.g., an estimate of 20 cm with a confidence interval of (14, 26)). While the standard errors are approximate (see Methods), our purpose is to approximately characterize uncertainty without delving into the details of alternative methods (e.g., the bootstrap) for estimating uncertainty.
@@
For many applications/locations, a 10560-member ensemble may be sufficient. However, if not, it is feasible now, and will become more so, to produce even larger ensembles, while cautioning that the benefit of this diminishes because relative statistical uncertainty decreases based on the inverse of the square root of the sample size.